A sample of platinum has a mass of 102g. When the sample is placed in a graduated cylinder containing 25.9mL of water, the water level rises to 30.3mL. What is the density of the platinum


1
Expert's answer
2021-09-23T04:51:54-0400

Volume of platinum "=30.3-25.9=4.4ml"


"Density =\frac{Mass}{Volume}"



"Density =\frac{102}{4.4}=23.18g/cm^3"